Hi, I need help with the following:
1. Solve by using substitution method. Determine whether equation is independent, dependent or inconsistent.
x=y+3
3x-2y=4

This is what I have so far.. Was I even on the right track?
X=y+3
X=y+3-3
X= -3

2. Solve by substitution or addition, whichever is easier
2y-x=3
X=3y-5

3. Solve by using addition method.
X+ y=13
22x+36y=356





Answer by checkley71(8403)

since x=(y+3) then we substitute (y+3) for x in the other equation thus
3(y+3)-2y=4
3y+9-2y=4
3y-2y=4-9
y=-5 solution now replace y with -5 & solve for x thus
x=(-5+3)
x=-2 solution
-----------------------------------------
x=3y-5 now substitute this in the other equation
2y-(3y-5)=3
2y-3y+5=3
-y=3-5
y=2 solution
x=3*2-5
x=6-5
x=1 solution
------------------------------------
x+y=13 then x=13-y
22(13-y)+36y=356
286-22y+36y=356
36y-22y=356-286
14y=70
y=70/14
y=5 solution
x+5=13
x=13-5
x=8 solution